# Symphony: A Heuristic Normalized Calibrated Advantage Actor and Critic Algorithm with Application to Humanoid Robots

### v6

@@ IV Experiments
Even with all this, there’s no $100\%$ guarantee that there will be absence of developmental defects. In our world, external obstacles often play an additional regulatory role. Following the logic of gradual human learning, we recommend starting training without obstacles, then gradually increasing the difficulty, filling a new Fading Replay Buffer with new exploratory data. While, for continuous learning with terminal cases to prevent overflow, we recommend subtracting a small value of $\epsilon$ from the ”done=1.0” feature, and then shifting it by 2 positions. This will ensure that after the second cycle, the terminal transition disappears unable to pass the next ”done equals 1.0” check.

### v7

@@ IV Experiments
Even with all this, there’s no $100\%$ guarantee that there will be absence of developmental defects. In our world, external obstacles often play an additional regulatory role. Following the logic of gradual human learning, we recommend starting training without obstacles, then gradually increasing the difficulty, filling a new Fading Replay Buffer with new exploratory data.
Symphony-$\tilde{S2}$
For practical use and for the upcoming ablation study, we decided to stick with Symphony-S2, with some modifications. For continuous learning with terminal cases to prevent overflow, we subtract a small value of $\epsilon$ from the ”done=1.0” feature, and then shifting it by 2 positions. If this value is below some critical threshold close to 1.0, we perform a standard shift by 1 position. This will ensure that after the N-th cycle, the terminal transition is overwritten. We additionally optimized the Fading Replay Buffer by refactoring the resource-consuming roll operation into a circular buffer with a cyclic pointer, while also implementing in-place operations to reduce memory overhead.
Instead of the modified hyperbolic tangent for the fixed weights, we used a bell-shaped Gaussian with a power of $2\pi$, a width of $1/\phi$, and a shift of $0.5+0.5\,(1/\phi)=\phi/2$, which has a maximum at the shift point $\phi/2$.
TABLE VI: Fixed weights
[Uncaptioned image] [Uncaptioned image]
a) for Target Critic b) for Fading Replay Buffer
$\omega_{i_{n}}=e^{-(\frac{|1-b-i_{n}|}{s})^{p}}$ $\omega_{i_{n}}=e^{-(\frac{|i_{n}-b|}{s})^{p}}$
$b=\phi/2$, $s=1/\phi$, $p=2\pi$
When using new weights for Target Critic, we reduce the importance of minimum values, while when using them for Fading Replay Buffer, our algorithm will not jump towards the latest transitions immediately.
Instead of 50% dropout probability or p=0.5, we used random probabilities (random values were generated using Gaussian Normal Distribution processed through Sigmoid function).
1) $x\leftarrow X\sim\mathcal{U}(0,1)$
2) $p\leftarrow sigmoid(P\sim\mathcal{N}(0,1))$
3) $mask\leftarrow\mathbb{I}(x>p)$